OpenStudy (anonymous):

Solve the system by elimination (What is the value of X, Y, & Z) -2x+y-2z=-10 -2x+y-z=-10 -x-y-3z=1

OpenStudy (anonymous):

solve the third quation for x x=-y-3z-1 then put the value of x on the othere equations -2(-y-3z-1)+y-2z=-10 2y+6z+2+y-2z=-10 3y+4z=-12 the second equation -2(-y-3z-1)+y-z=-10 2y+6z+2+y-z=-10 3y+5z=-12 now sloving 3y+5z=-12 for y 3y=(-12-5z) putting this value for y in the othere equation.. 3y+4z=-12 -12-5z+4z=-12 -z=-12+12=0 solving for 3y+4z=-12...... 3y+4(0)=-12.... y=-12/3=-4 soving for x -2x+y-z=-10 x=(-10+4)/-2=-6/-2=3